# FedDyMem: Efficient Federated Learning with Dynamic Memory and Memory-Reduce for Unsupervised Image Anomaly Detection

### v1

@@ VI Experiments
VI-A Experimental Settings
VI-A1 Datasets
In this article, we combined $6$ image anomaly detection datasets containing different types of data, covering various industrial inspection scenarios and medical imaging domains, to evaluate the performance of FedDyMem. These datasets are as follows:
  • Mixed-Brain-AD(Brain): We combined $3$ brain image diagnostic datasets, including Brain-Tumor [2], BraTS2021  [^1] https://www.kaggle.com/datasets/masoudnickparvar/brain-tumor-mri-dataset [68] and HeadCT  [^2] https://www.kaggle.com/datasets/felipekitamura/head-ct-hemorrhage [69]. These datasets, consisting of both MRI and CT images, were distributed across different clients in a heterogeneous manner to represent a mixture of various image types.
  • Mixed-ChestXray-AD(X-ray): Three diagnostic chest X-ray datasets, including ChestXRay2017[70], RSNA [^3] https://www.kaggle.com/c/rsna-pneumonia-detection-challenge[71], and VinDr-CXR [^4] https://www.kaggle.com/c/vinbigdata-chest-xray-abnormalities-detection[72], were employed in this article. These datasets were derived from distinct medical institutions, ensuring a diverse and heterogeneous data distribution, which is essential for the federated UAD.
  • Mixed-PCB-AD(PCB): We investigate $2$ datasets containing PCB anomalies, namely VisA[73] and VISION[74] , and integrate them to create the Mixed-PCB dataset. The VisA dataset includes six different types of PCB products, while the VISION dataset encompasses two types. Consequently, the Mixed-PCB-AD dataset provides a total of $8$ distinct types of PCB product images, facilitating simulation in federated UAD.
  • MVTec[75]: In this article, we utilize $5$ texture classes(carpet, grid, leather, tile, and wood) from the MVTec dataset, which includes mask annotations, to simulate industrial federated anomaly detection scenarios.
  • MetalPartsAD(MPDD)[76]: The MPDD comprises images of $6$ types of metal parts, each captured under varying conditions of spatial orientation, object positioning, and camera distance. These images are further characterized by diverse lighting intensities and non-homogeneous backgrounds, providing a comprehensive dataset for our federated UAD.
  • TextureAD-Wafer(Wafer)[19]: TextureAD-Wafer comprises a dataset of 14 distinct wafer products, each imaged using an industrial-grade, high-resolution optical camera. This dataset also explores these wafer products under various optical conditions, thus providing a rich federated UAD simulate environment with diverse feature distributions.
@@ VI-A1 Datasets
Figure 6: Illustration of the loss curves across five clients evaluated on four datasets, demonstrating convergence trends and performance variability within our experiments.
VI-A2 Scenes with Feature Distribution Bias
@@
In this article, we aim to simulate heterogeneous feature distributions across $5$ clients by all samples from the same normal categories with different product types. We employ a Dirichlet distribution to allocate data, where each product type, though labeled as normal, is distributed to various clients. We set the Dirichlet parameter $\alpha=0.1$ to accentuate the variances in data distribution across clients, thus simulating more realistic and challenging multi-client scenarios typically encountered in federated UAD. Fig. 5 illustrates the results of this partitioning approach across all datasets.
VI-A3 Implementation Details
@@
The frozen feature extractor used in our experiments is the Wide-ResNet-50 model pre-trained on the ImageNet dataset. For FedDyMem, the hyperparameters $H^{\bf G}$ and $W^{\cal G}$ are set to $8$, and $th$ is set to $0.01$. All baseline models are constructed using the Wide-ResNet-50, with all parameters comprehensively aggregated. The parameter $K$ for the KNN algorithm is set to 3. For a fair comparison with other state-of-the-art methods, we ensure consistency in the training parameters used in our experiments. Specifically, the learning rate is set to $1\times 10^{-3}$, with a total of 200 communication rounds. The local epoch is set to 1, and the batch size is 10. We utilize the Adam optimizer with a weight decay of $5\times 10^{-4}$ and a momentum of $1$. All images in the dataset were preprocessed by cropping and resizing to a resolution of $224\times 224$ pixels using the bicubic interpolation method.
VI-A4 Evaluation Metric
@@
For the evaluation of anomaly detection, we adopt the Area Under the Receiver Operating Characteristic Curve (AUROC), ensuring a comprehensive assessment of our proposed model in alignment with prior studies. Image-level anomaly detection performance is quantified using the standard AUROC metric, denoted as Image-level AUROC(I-AUROC). Additionally, to assess anomaly localization, we compute the Pixel-level AUROC(P-AUROC). To further enhance the evaluation of anomaly localization capabilities, we calculate the Per-Region Overlap (PRO) at the pixel level, providing a more granular and comprehensive analysis. These metrics are widely used in previous work [51, 36, 17, 2].
TABLE I: Comparison of Image-AUROC (%) Results on Six Dataset
Methods Communication Parameters Wafer PCB MVTec MPDD Brain X-ray
@@
FedAvg Model Paramters: $\approx$ 10.66M, (SCAFFOLD) $\approx$ 10.66M * 2 63.91($\pm$0.36) 54.18($\pm$0.33) 80.32($\pm$0.81) 61.87($\pm$0.76) 50.16($\pm$0.63) 61.48($\pm$0.56)
@@
VI-B Convergence Curves
In this section, we present a detailed analysis of the training loss curves for FedDyMem in comparison with baseline models across four datasets including Brain, Wafer, PCB, and X-ray. The loss curves for all $5$ clients are illustrated in Fig. 6 over $200$ communication rounds, providing insights into the performance and convergence behavior of the models. First, FedDyMem demonstrates a significantly faster convergence rate across all datasets, achieving near-convergence within only a few communication rounds. This rapid convergence shows the efficiency and robustness of FedDyMem in distributed learning scenarios, making it particularly suitable for real-world applications where communication efficiency is critical. Second, FedDyMem exhibits smoother convergence compared to the baseline models. As shown in the zoomed-in view of the later communication rounds in Fig. 6, the loss curves of the baseline models have less stable optimization dynamics. In contrast, FedDyMem maintains consistent and stable convergence, highlighting its superior optimization process. Moreover, FedDyMem consistently achieves lower final loss values across all datasets, outperforming baseline models in terms of both accuracy and stability. This superior performance can be attributed to FedDyMem’s dynamic memory bank mechanism, which facilitates superior global generalization. The experimental results substantiate the convergence analysis presented in Section V, further validating the efficiency of FedDyMem in federated learning settings.
VI-C Performance Comparisons with Existing Methods
To demonstrate the superior performance of FedDyMem, we conducted extensive comparative analyses with $8$ baseline models. These include the basic federated learning algorithm FedAvg [54], the regularization-based FedProx[22], the personalized-based FedPHP[77] and APPLE[29]. Additionally, we compared against domain-skew-oriented methods such as FedBN [23] and SCAFFOLD [20], as well as the model-splitting method DBE [35] and a purely local training scheme (Local).
TABLE II: Comparison of Abnormal Location Capability on Four Datasets
@@ VI-C Performance Comparisons with Existing Methods
P-AUROC(%) PRO(%) P-AUROC(%) PRO(%) P-AUROC(%) PRO(%) P-AUROC(%) PRO(%)
Local 58.27($\pm$1.20) 47.76($\pm$0.52) 68.12($\pm$1.11) 45.86($\pm$1.31) 72.63($\pm$0.98) 56.97($\pm$0.48) 87.02($\pm$1.05) 51.31($\pm$0.73)
FedAvg 66.56($\pm$0.70) 67.00($\pm$0.30) 69.30($\pm$0.34) 54.04($\pm$0.57) 76.16($\pm$0.73) 53.35($\pm$1.49) 91.67($\pm$0.94) 55.07($\pm$1.25)
FedProx 66.41($\pm$0.74) 67.25($\pm$0.33) 69.10($\pm$1.28) 54.04($\pm$1.28) 76.08($\pm$1.28) 53.62($\pm$0.70) 91.59($\pm$1.02) 54.96($\pm$1.16)
FedPHP 66.44($\pm$1.38) 67.16($\pm$0.38) 69.11($\pm$0.42) 54.23($\pm$1.16) 76.03($\pm$0.57) 53.37($\pm$0.47) 91.56($\pm$0.73) 54.86($\pm$1.42)
APPLE 69.88($\pm$0.95) 69.02($\pm$0.42) 71.34($\pm$0.85) 60.11($\pm$1.45) 78.06($\pm$1.08) 54.14($\pm$1.16) 92.12($\pm$1.19) 55.84($\pm$0.76)
FedBN 66.32($\pm$0.57) 67.16($\pm$0.68) 68.91($\pm$0.82) 54.21($\pm$1.07) 75.92($\pm$1.34) 53.48($\pm$0.79) 91.59($\pm$0.38) 54.92($\pm$0.82)
SCAFFOLD 68.75($\pm$1.19) 68.74($\pm$1.03) 77.18($\pm$0.54) 66.43($\pm$0.58) 80.81($\pm$0.35) 55.79($\pm$1.30) 93.47($\pm$0.74) 58.71($\pm$0.84)
DBE 66.95($\pm$1.07) 68.46($\pm$1.05) 72.95($\pm$1.35) 57.89($\pm$0.75) 76.18($\pm$0.89) 53.86($\pm$0.58) 92.66($\pm$1.36) 55.76($\pm$0.95)
Ours 78.40($\pm$1.31) 74.91($\pm$0.54) 96.03($\pm$0.79) 81.09($\pm$1.31) 97.04($\pm$0.56) 86.81($\pm$1.13) 98.97($\pm$1.39) 79.43($\pm$1.15)
VI-C1 Training Process
@@
Fig. 7 illustrates the I-AUROC performance of the FedDyMem training process in comparison with baseline models on all datasets. FedDyMem achieves superior testing accuracy compared to other baseline models, while also exhibiting a more stable training process. Specifically, the baseline models exchange parameters exclusively during communication rounds, introducing a distribution bias for memory features. This limitation prevents the models from achieving suboptimal solutions.
VI-C2 Communication Cost
@@
In federated learning, communication efficiency is crucial for practical deployment, especially in bandwidth-constrained or latency-sensitive environments. As shown in Table I, most existing methods (e.g., FedAvg, FedProx, APPLE) require extensive communication of full model parameters, which amount to approximately 10.66 MB per communication round. In contrast, our approach only requires the exchange of a memory bank of approximately 5.62MB during each communication round. This substantially reduces communication overhead, enabling faster convergence in settings with limited communication resources.
VI-C3 Anomaly Detection
@@
As shown in Table I, the performance of FedDyMem is evaluated across six datasets and compared with a set of baseline models. Conventional federated learning methods such as FedAvg, FedProx, and FedPHP show similar performance, typically outperforming the Local, but remaining below 65% AUROC on most datasets (e.g., 64% on Wafer). APPLE and FedBN show improvements over FedAvg-like methods (e.g., 65.92% AUROC on Wafer for APPLE), but their effectiveness decreases on more challenging datasets such as Brain, where AUROC remains below 51%. SCAFFOLD achieves relatively better results (e.g., 72.52% on X-ray), but provides limited gains on PCB, Brain, and MPDD. Similarly, DBE shows incremental improvements across all benchmarks but fails to match our method. In contrast, our proposed approach consistently achieves the highest I-AUROC scores across all six anomaly detection benchmarks. Specifically, our method achieves 71.96% for Wafer, 83.34% for PCB, 99.24% for MVTec, 83.93% for MPDD, 89.27% for Brain, and 80.30% for X-ray.
VI-C4 Anomaly Localization
@@
As shown in Table II we evaluate the abnormal localization performance on four datasets (Wafer, PCB, MVtec, and MPDD), as the Brain and X-ray do not provide region-level labels. To measure both detection capability and localization accuracy, each method is evaluated using two metrics, P-AUROC and PRO. The baseline method Local exhibits the lowest overall performance, with approximately 58.27% in P-AUROC and 47.76% in PRO on the Wafer dataset, and similar results across the other datasets. This proves that there are distribution bias in the raw dataset. For the federated learning approaches, FedAvg, FedProx, and FedPHP demonstrate closely aligned performance improvements. APPLE delivers further performance gains, achieving a P-AUROC of up to 69.88% on the Wafer dataset. In contrast, our proposed method achieves superior performance across all evaluated datasets and metrics. These consistent and robust improvements highlight the effectiveness and adaptability of our approach in tackling federated anomaly localization tasks across diverse application scenarios.
TABLE III: Ablation Study of Structure for FedDyMem.
+Projection +Memory Generator +$K$-means I-AUROC (%) P-AUROC (%)
@@
VI-D Ablation Study
VI-D1 Impact of Local Training Epochs
Fig. 8 illustrates the impact of varying the number of local training epochs ($E$) on the performance of federated learning across all datasets. A clear trend emerges where increasing $E$ generally results in a performance decline. For example, in the Wafer dataset, accuracy drops from 71.96% at $E=1$ to 69.92% at $E=5$. Similarly, the PCB dataset demonstrates a decrease in performance, with accuracy falling from 83.34% at $E=1$ to 82.65% at $E=5$. MPDD, Brain, and X-ray also exhibit gradual performance degradation with higher $E$ values. This phenomenon can be attributed to the heterogeneous data distribution across clients. As the number of local training epochs increases, the local models tend to overfit their respective client datasets, leading to a decrease in the effectiveness of the globally aggregated memory features.

### v2

@@ 6 Experiments
6.1 Experimental Settings
Datasets. In this article, we combined $6$ image anomaly detection datasets containing different types of data, covering various industrial inspection scenarios and medical imaging domains, to evaluate the performance of FedDyMem. These datasets are as follows:
  • Mixed-Brain-AD(Brain): We combined $3$ brain image diagnostic datasets, including Brain-Tumor ([cai2024medianomaly]), BraTS2021 ([baid2021rsna]) and HeadCT ([salehi2021multiresolution]). These datasets, consisting of both MRI and CT images, were distributed across different clients in a heterogeneous manner to represent a mixture of various image types.
  • Mixed-ChestXray-AD(X-ray): Three diagnostic chest X-ray datasets, including ChestXRay2017 ([kermany2018identifying]), RSNA ([wang2017chestx]), and VinDr-CXR ([nguyen2022vindr]), were employed in this article. These datasets were derived from distinct medical institutions, ensuring a diverse and heterogeneous data distribution, which is essential for the federated UAD.
  • Mixed-PCB-AD(PCB): We investigate $2$ datasets containing PCB anomalies, namely VisA ([zou2022spot]) and VISION ([bai2023vision]), and integrate them to create the Mixed-PCB dataset. The VisA dataset includes six different types of PCB products, while the VISION dataset encompasses two types. Consequently, the Mixed-PCB-AD dataset provides a total of $8$ distinct types of PCB product images, facilitating simulation in federated UAD.
  • MVTec ([bergmann2019mvtec]): In this article, we utilize $5$ texture classes(carpet, grid, leather, tile, and wood) from the MVTec dataset, which includes mask annotations, to simulate industrial federated anomaly detection scenarios.
  • MetalPartsAD(MPDD) ([jezek2021deep]): The MPDD comprises images of $6$ types of metal parts, each captured under varying conditions of spatial orientation, object positioning, and camera distance. These images are further characterized by diverse lighting intensities and non-homogeneous backgrounds, providing a comprehensive dataset for our federated UAD.
  • TextureAD-Wafer(Wafer) ([lei2025adapted]): TextureAD-Wafer comprises a dataset of 14 distinct wafer products, each imaged using an industrial-grade, high-resolution optical camera. This dataset also explores these wafer products under various optical conditions, thus providing a rich federated UAD simulate environment with diverse feature distributions.
@@ 6.1 Experimental Settings
Scenes with Feature Distribution Bias. In this article, we aim to simulate heterogeneous feature distributions across $5$ clients by using all samples from the same normal categories with different product types. We employ a Dirichlet distribution to allocate data, where each product type, though labeled as normal, is distributed to various clients. We set the Dirichlet parameter $\alpha=0.1$ to accentuate the variances in data distribution across clients, thus simulating more realistic and challenging multi-client scenarios typically encountered in federated UAD. Fig. 5 illustrates the results of this partitioning approach across all datasets.
@@
Implementation Details. The frozen feature extractor used in our experiments is the Wide-ResNet-50 model pre-trained on the ImageNet dataset. For FedDyMem, the hyperparameters $H^{\bf G}$ and $W^{\cal G}$ are set to $8$, and $th$ is set to $0.01$. All baseline models are constructed using the Wide-ResNet-50, with all parameters comprehensively aggregated. The parameter $K$ for the KNN algorithm is set to 3. For a fair comparison with other state-of-the-art methods, we ensure consistency in the training parameters used in our experiments. Specifically, the learning rate is set to $1\times 10^{-3}$, with a total of 200 communication rounds. The local epoch is set to 1, and the batch size is 10. We utilize the Adam optimizer with a weight decay of $5\times 10^{-4}$ and a momentum of $1$. All images in the dataset were preprocessed by cropping and resizing to a resolution of $224\times 224$ pixels using the bicubic interpolation method.
@@
Evaluation Metric. For the evaluation of anomaly detection, we adopt the Area Under the Receiver Operating Characteristic Curve (AUROC), ensuring a comprehensive assessment of our proposed model in alignment with prior studies. Image-level anomaly detection performance is quantified using the standard AUROC metric, denoted as Image-level AUROC(I-AUROC). Additionally, to assess anomaly localization, we compute the Pixel-level AUROC(P-AUROC). To further enhance the evaluation of anomaly localization capabilities, we calculate the Per-Region Overlap (PRO) at the pixel level, providing a more granular and comprehensive analysis. These metrics are widely used in previous work ([roth2022towards, you2022unified, liang2023omni, cai2024medianomaly]).
Figure 6: Illustration of the loss curves across five clients evaluated on four datasets. To better highlight the effectiveness of FedDyMem, zoom-in views of the later communication rounds are included in each subfigure. In the zoom-in, we annotate three key values: (1) the loss of FedDyMem (Ours) as the first row, (2) the maximum loss among all compared methods in the second row, and (3) the minimum loss excluding our method in the third row.
Figure 7: Comparison of feature distributions on the MVTec dataset after 200 communication rounds. (a) t-SNE ([maaten2008visualizing]) visualization of local and global memory banks generated by five clients using FedAvg (top row) and FedDyMem (bottom row). (b) Comparison of global memory banks aggregated from FedDyMem and FedAvg clients, together with the corresponding I-AUROC results.
6.2 Convergence Curves and Distribution Consistency
In this section, we present a detailed analysis of the training loss curves for FedDyMem in comparison with baseline models across four datasets including Brain, Wafer, PCB, and X-ray. The loss curves for all $5$ clients are illustrated in Fig. 6 over $200$ communication rounds, providing insights into the performance and convergence behavior of the models. First, FedDyMem demonstrates a significantly faster convergence rate across all datasets. Second, FedDyMem exhibits smoother convergence compared to the baseline models. As shown in the zoomed-in view of the later communication rounds in Fig. 6, the loss curves of the baseline models have less stable optimization dynamics. In contrast, FedDyMem maintains consistent and stable convergence. Moreover, FedDyMem consistently achieves lower final loss values across all datasets. The experimental results substantiate the convergence analysis presented in Section 5, further validating the efficiency of FedDyMem in federated learning settings.
To further analyze the discrepancy between the loss values in Fig. 6 and the AUROC improvements reported in Table 1, we visualize in Fig. 7 the t-SNE of the memory feature distributions learned by FedAvg and FedDyMem on the MVTec dataset after 200 communication rounds. Specifically, we compare the local memory banks of five clients against their corresponding global memory banks in Fig. 7(a). For FedAvg, the memory features generated by individual clients exhibit clear distributional deviations from the global memory, leading to local losses that remain several orders of magnitude higher (up to $10^{4}$) than those of FedDyMem. This mismatch indicates that although the optimization objective in FedAvg continues to decrease, the learned memory features fail to align with the global feature distribution due to local feature bias and inter-client distribution shifts. Furthermore, Fig. 7(b) illustrates that FedDyMem achieves a much more compact and coherent global feature space across clients. The memory features aggregated via dynamic memory update and k-means clustering form a well-aligned global representation that mitigates the impact of intra-class distribution bias. This alignment is critical for UAD, where the effectiveness of nearest-neighbor-based scoring relies heavily on the quality and consistency of the global memory. Therefore, the superior AUROC performance of FedDyMem benefits from its ability to construct globally consistent and discriminative feature spaces across heterogeneous clients.
Table 1: Comparison of Image-AUROC (%) results on six Dataset
Methods Communication Wafer PCB MVTec MPDD Brain X-ray
@@ 6.2 Convergence Curves and Distribution Consistency
FedAvg Model Parameters: $\approx$ 10.66M, (SCAFFOLD) $\approx$ 10.66M * 2 63.91 ($\pm$0.36) 54.18 ($\pm$0.33) 80.32 ($\pm$0.81) 61.87 ($\pm$0.76) 50.16 ($\pm$0.63) 61.48 ($\pm$0.56)
@@
6.3 Performance Comparisons with Existing Methods
To demonstrate the superior performance of FedDyMem, we conducted extensive comparative analyses with $8$ baseline models. These include the basic federated learning algorithm FedAvg ([fedavg]), the regularization-based FedProx ([fedprox]), the personalized-based FedPHP ([li2021fedphp]) and APPLE ([luo2022adapt]). Additionally, we compared against domain-skew-oriented methods such as FedBN ([li2021fedbn]) and SCAFFOLD ([karimireddy2020scaffold]), as well as the model-splitting method DBE ([feddbe]) and a purely local training scheme (Local).
Table 2: Comparison of Pixel-AUROC (%) results on four Dataset
@@ 6.3 Performance Comparisons with Existing Methods
Local 58.27($\pm$1.20) 68.12($\pm$1.11) 72.63($\pm$0.98) 87.02($\pm$1.05)
FedAvg 66.56($\pm$0.70) 69.30($\pm$0.34) 76.16($\pm$0.73) 91.67($\pm$0.94)
FedProx 66.41($\pm$0.74) 69.10($\pm$1.28) 76.08($\pm$1.28) 91.59($\pm$1.02)
FedPHP 66.44($\pm$1.38) 69.11($\pm$0.42) 76.03($\pm$0.57) 91.56($\pm$0.73)
APPLE 69.88($\pm$0.95) 71.34($\pm$0.85) 78.06($\pm$1.08) 92.12($\pm$1.19)
FedBN 66.32($\pm$0.57) 68.91($\pm$0.82) 75.92($\pm$1.34) 91.59($\pm$0.38)
SCAFFOLD 68.75($\pm$1.19) 77.18($\pm$0.54) 80.81($\pm$0.35) 93.47($\pm$0.74)
DBE 66.95($\pm$1.07) 72.95($\pm$1.35) 76.18($\pm$0.89) 92.66($\pm$1.36)
Ours 78.40($\pm$1.31) 96.03($\pm$0.79) 97.04($\pm$0.56) 98.97($\pm$1.39)
Table 3: Comparison of PRO (%) results on four Dataset
Methods Wafer PCB MVTec MPDD
Local 47.76($\pm$0.52) 45.86($\pm$1.31) 56.97($\pm$0.48) 51.31($\pm$0.73)
FedAvg 67.00($\pm$0.30) 54.04($\pm$0.57) 53.35($\pm$1.49) 55.07($\pm$1.25)
FedProx 67.25($\pm$0.33) 54.04($\pm$1.28) 53.62($\pm$0.70) 54.96($\pm$1.16)
FedPHP 67.16($\pm$0.38) 54.23($\pm$1.16) 53.37($\pm$0.47) 54.86($\pm$1.42)
APPLE 69.02($\pm$0.42) 60.11($\pm$1.45) 54.14($\pm$1.16) 55.84($\pm$0.76)
FedBN 67.16($\pm$0.68) 54.21($\pm$1.07) 53.48($\pm$0.79) 54.92($\pm$0.82)
SCAFFOLD 68.74($\pm$1.03) 66.43($\pm$0.58) 55.79($\pm$1.30) 58.71($\pm$0.84)
DBE 68.46($\pm$1.05) 57.89($\pm$0.75) 53.86($\pm$0.58) 55.76($\pm$0.95)
Ours 74.91($\pm$0.54) 81.09($\pm$1.31) 86.81($\pm$1.13) 79.43($\pm$1.15)
@@
Training Process. Fig. 8 illustrates the I-AUROC performance of the FedDyMem training process in comparison with baseline models on all datasets. FedDyMem achieves superior testing accuracy compared to other baseline models, while also exhibiting a more stable training process. Specifically, the baseline models exchange parameters exclusively during communication rounds, introducing a distribution bias for memory features. This limitation prevents the models from achieving suboptimal solutions.
@@
Communication Cost. In federated learning, communication efficiency is crucial for practical deployment, especially in bandwidth-constrained or latency-sensitive environments. As shown in Table 1, most existing methods (e.g., FedAvg, FedProx, APPLE) require extensive communication of full model parameters, which amount to approximately 10.66 MB per communication round. In contrast, our approach only requires the exchange of a memory bank of approximately 5.62MB during each communication round. This substantially reduces communication overhead, enabling faster convergence in settings with limited communication resources.
@@
Anomaly Detection. As shown in Table 1, the performance of FedDyMem is evaluated across six datasets and compared with a set of baseline models. Conventional federated learning methods such as FedAvg, FedProx, and FedPHP show similar performance, typically outperforming the Local, but remaining below 65% AUROC on most datasets (e.g., 64% on Wafer). APPLE and FedBN show improvements over FedAvg-like methods (e.g., 65.92% AUROC on Wafer for APPLE), but their effectiveness decreases on more challenging datasets such as Brain, where AUROC remains below 51%. SCAFFOLD achieves relatively better results (e.g., 72.52% on X-ray), but provides limited gains on PCB, Brain, and MPDD. Similarly, DBE shows incremental improvements across all benchmarks but fails to match our method. In contrast, our proposed approach consistently achieves the highest I-AUROC scores across all six anomaly detection benchmarks. Specifically, our method achieves 71.96% for Wafer, 83.34% for PCB, 99.24% for MVTec, 83.93% for MPDD, 89.27% for Brain, and 80.30% for X-ray.
@@
Anomaly Localization. As shown in Table 2 and Table 3, we evaluate the abnormal localization performance on four datasets (Wafer, PCB, MVtec, and MPDD), as the Brain and X-ray do not provide region-level labels. To measure both detection capability and localization accuracy, each method is evaluated using two metrics, P-AUROC and PRO. The baseline method Local exhibits the lowest overall performance, with approximately 58.27% in P-AUROC and 47.76% in PRO on the Wafer dataset, and similar results across the other datasets. This proves that there is distribution bias in the raw dataset. For the federated learning approaches, FedAvg, FedProx, and FedPHP demonstrate closely aligned performance improvements. APPLE delivers further performance gains, achieving a P-AUROC of up to 69.88% on the Wafer dataset. In contrast, our proposed method achieves superior performance across all evaluated datasets and metrics. These consistent and robust improvements highlight the effectiveness and adaptability of our approach in tackling federated anomaly localization tasks across diverse application scenarios.
Table 4: Ablation study of structure for FedDyMem.
+Proj +MG +KM I-AUROC (%) P-AUROC (%)
@@
6.4 Ablation Study
Impact of Local Training Epochs. Fig. 9 illustrates the impact of varying the number of local training epochs ($E$) on the performance of federated learning across all datasets. A clear trend emerges where increasing $E$ generally results in a performance decline. For example, in the Wafer dataset, accuracy drops from 71.96% at $E=1$ to 69.92% at $E=5$. Similarly, the PCB dataset demonstrates a decrease in performance, with accuracy falling from 83.34% at $E=1$ to 82.65% at $E=5$. MPDD, Brain, and X-ray also exhibit gradual performance degradation with higher $E$ values. This phenomenon can be attributed to the heterogeneous data distribution across clients. As the number of local training epochs increases, the local models tend to overfit their respective client datasets, leading to a decrease in the effectiveness of the globally aggregated memory features.
Effect of Projection Layer. The ablation study presented in Table 4 shows the significant impact of incorporating a projection layer (Proj) on the performance of the proposed FedDyMem framework. The baseline configuration used for comparison consists of memory-reduction and memory feature average aggregation mechanisms, without additional architectural enhancements. Specifically, introducing the projection layer improves both I-AUROC and P-AUROC metrics across all datasets. For example, the I-AUROC for Wafer increases from 67.69% to 68.93%, and the P-AUROC improves from 70.23% to 74.29%. These results demonstrate that the projection layer effectively enhances feature representations.
Effect of Memory Generator. The integration of the memory generator (MG) significantly improves the anomaly detection performance of FedDyMem across different datasets. For example, the memory generator increases the I-AUROC from 68.93% to 70.96% on the wafer dataset and from 74.10% to 81.26% on the PCB dataset. A similar trend is observed for P-AUROC, with increases from 74.29% to 77.01% on Wafer and from 89.26% to 94.97% on PCB. These improvements demonstrate the ability of the memory generator to effectively refine feature representations.
Effect of $K$-means. The addition of $K$-means (KM) to the FedDyMem framework resulted in a consistent improvement across all evaluated metrics. When combined with the projection layer and memory generator, $K$-means further improves I-AUROC, achieving increases such as from 70.96% to 71.96% for Wafer and from 81.26% to 83.34% for PCB. Similarly, P-AUROC improves significantly with values such as Wafer from 77.01% to 78.40% and PCB from 94.97% to 96.03%.